• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 00:35:03 on November 7,2006

    Entry number 189023

    This entry is a followup to: 189016

    keyword=poor sig/bgr, stop Compton tune

    Alex tried varying the beam current to find a good signal-to-bgr
    for Compton. Last week when the current was low this ratio improved
    dramatically (it was not understood why). But tonight it actually
    got worse with low current. Alex also tried the dogleg adjustment.
    No improvement. We are stopping now. Back to 85 uA production.
    There is basically no signal in Compton, sig:bgr = 1:20 approx.